# Exploiting Completeness Perception with Diffusion Transformer for Unified 3D MRI Synthesis

### v1

@@ 3.2 Stage I: Completeness Perception Tokenizer
LDM [39] 23.84 $\pm$1.52 0.805 $\pm$0.019 36.47 $\pm$1.54 23.12 $\pm$1.64 0.798 $\pm$0.021 45.93 $\pm$2.41 22.65 $\pm$1.77 0.791 $\pm$0.025 52.50 $\pm$2.07 21.36 $\pm$1.28 0.679 $\pm$0.016 86.62 $\pm$3.86 20.94 $\pm$1.53 0.654 $\pm$0.020 112.57 $\pm$5.02
ControlNet [58] 23.98 $\pm$1.49 0.808 $\pm$0.018 34.09 $\pm$1.93 23.34 $\pm$1.60 0.801 $\pm$0.020 41.28 $\pm$2.49 22.85 $\pm$1.68 0.795 $\pm$0.022 48.07 $\pm$1.99 21.83 $\pm$1.51 0.681 $\pm$0.015 81.26 $\pm$3.85 21.07 $\pm$1.50 0.661 $\pm$0.019 103.77 $\pm$5.17
M2DN [29] 26.45 $\pm$1.36 0.830 $\pm$0.016 21.29 $\pm$1.18 25.87 $\pm$1.48 0.820 $\pm$0.017 27.36 $\pm$1.62 25.21 $\pm$1.59 0.809 $\pm$0.024 32.40 $\pm$1.93 23.47 $\pm$1.43 0.715 $\pm$0.014 42.52 $\pm$3.08 22.81 $\pm$1.56 0.702 $\pm$0.018 55.64 $\pm$3.62
CoPeDiT 28.26 $\pm$1.24 0.842 $\pm$0.019 12.67 $\pm$0.98 28.13 $\pm$1.49 0.831 $\pm$0.021 13.25 $\pm$1.45 27.91 $\pm$1.41 0.822 $\pm$0.023 14.89 $\pm$1.61 24.34 $\pm$1.21 0.732 $\pm$0.016 25.84 $\pm$2.58 23.92 $\pm$1.53 0.721 $\pm$0.020 32.53 $\pm$3.16
@@ 3.3 Stage II: 3D MRI Diffusion Transformer
We propose MDiT3D (see Fig. 3), extending the model configuration of DiT [34] with several key design innovations to enable MDiT3D to adapt to 3D modalities and volumetric MRI. It takes the high-dimensional latent tokens $\mathbf{z}^{s}$ as input, along with the learned prompt tokens $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$ from our CoPeVAE tokenizer, which encode comprehensive missing state information and serve as conditional guidance for synthesizing. Without loss of generality, we use MDiT3D-B as a representative configuration to illustrate the architectural design.
3D Patchify & 3D Positional Embeddings. We apply a 3D patchify operator with patch size $p\times p\times p$ to obtain $\mathbf{h}\in\mathbb{R}^{m\times T\times d}$ [30], where $m$ denotes the number of modality, and $d$ is the embedding dimension. Unlike the standard 2D positional embeddings [9] used in DiT, we leverage 3D frequency-based sine-cosine positional embeddings (3D PE) to better encode spatial positional information within the 3D MRI data.
Alternating Blocks & Prompt Injection. Two distinct types of blocks are designed tailored for brain and cardiac tasks, respectively. For the brain MRI synthesis task, we adopt spatial and modal blocks: the former focuses on capturing 3D spatial information, while the latter models inter-modal relationships. Before each modal block, the tokens $\mathbf{h}$ are reshaped into $\mathbb{R}^{T\times m\times d}$ to facilitate inter-modal interaction, then reshaped back for subsequent spatial processing. These two types of blocks are stacked in an alternating fashion to form a total of $N$ blocks. The motivation behind this design lies in the need for the model to capture both inter- and intra-modal relationships to boost contextual understanding and anatomical consistency. Additionally, adaptive layer normalization is employed to inject prompts into the blocks. Notably, inspired by the intuition that conditioning signals should be subtle yet informative, prompt tokens are injected only into the modal blocks.
Joint Reconstruction & Synthesis. We concatenate the available and the noisy latent (corresponding to missing sections) as the input to MDiT3D. Moreover, a joint reconstruction & synthesis scheme is introduced, where only the missing sections undergo noise injection and denoising during the diffusion process, while the observed regions remain unchanged, serving as auxiliary context that provides rich semantic guidance for generation. Given that directly synthesizing the entire modality/volume is more informative and efficient than simply predicting zero noise for the available latent regions [29], we adopt x-prediction loss as objective function for MDiT3D, as below:
@@
$\small\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{z}_{0},\epsilon,t}\left[\left\|\mathbf{z}_{0}-f_{\theta}(\mathbf{z}_{t},t,\mathbf{p})\right\|_{2}^{2}\right].$ (5)
@@
where $\mathbf{z}_{0}=\mathbf{z}^{s}\,\|\,\mathbf{z}^{c}$ denotes the concatenated clean latent tokens, $\mathbf{z}_{t}=\mathbf{z}^{s}\,\|\,\mathbf{z}_{t}^{c}$ is the noisy input at timestep $t$, $\mathbf{p}$ is the prompt tokens, and $f_{\theta}$ represents our MDiT3D model.
4 Experiments and Results
@@
Table 2: Quantitative results for Cardiac MRI synthesis on UKBB dataset. The numbers of the second row denote the length of missing slices. The best performances are bolded.
UKBB
@@
MMGAN [43] 25.81 $\pm$0.86 0.815 $\pm$0.012 16.68 $\pm$0.93 24.35 $\pm$0.84 0.793 $\pm$0.009 27.52 $\pm$0.90 23.06 $\pm$1.08 0.776 $\pm$0.018 45.88 $\pm$1.32
MMT [24] 26.02 $\pm$0.82 0.824 $\pm$0.009 15.90 $\pm$0.88 24.73 $\pm$0.87 0.809 $\pm$0.010 23.75 $\pm$0.97 24.12 $\pm$1.06 0.794 $\pm$0.017 37.39 $\pm$1.27
Hyper-GAE [53] 25.23 $\pm$0.89 0.810 $\pm$0.012 19.02 $\pm$0.96 23.66 $\pm$0.94 0.789 $\pm$0.013 31.84 $\pm$1.15 22.70 $\pm$1.14 0.771 $\pm$0.019 48.27 $\pm$1.39
@@
LDM [39] 24.17 $\pm$0.96 0.795 $\pm$0.012 24.61 $\pm$0.94 23.04 $\pm$1.03 0.778 $\pm$0.010 44.93 $\pm$1.49 22.19 $\pm$1.21 0.761 $\pm$0.018 60.02 $\pm$1.68
ControlNet [58] 24.62 $\pm$0.91 0.801 $\pm$0.012 22.47 $\pm$1.06 23.30 $\pm$0.98 0.784 $\pm$0.014 37.51 $\pm$1.55 22.46 $\pm$1.16 0.765 $\pm$0.020 54.93 $\pm$1.74
M2DN [29] 25.48 $\pm$0.79 0.814 $\pm$0.011 17.72 $\pm$0.85 24.62 $\pm$0.87 0.803 $\pm$0.011 24.15 $\pm$1.06 24.03 $\pm$0.99 0.780 $\pm$0.021 40.08 $\pm$1.59
CoPeDiT 26.42 $\pm$0.81 0.831 $\pm$0.013 15.53 $\pm$0.88 26.07 $\pm$0.74 0.826 $\pm$0.013 18.21 $\pm$0.90 25.39 $\pm$0.86 0.817 $\pm$0.016 25.84 $\pm$1.22
Figure 4: Qualitative results on BraTS dataset. The results are depicted in the axial (top), sagittal (middle) and coronal (bottom) views of the 3D MRI volume. The visual results on IXI dataset are provided in the supplementary material.
Figure 5: Qualitative results of comparison on UKBB dataset. The top and bottom results correspond to the first and last missing slices within a given volume, respectively.
Table 3: Ablation study on the contribution of each pretext task to CoPeVAE’s reconstruction capacity. ‘Cardiac’ refers to the aggregated cardiac MRI dataset consisting of four sources, which are used for training our CoPeVAE model.
@@
Table 4: Ablation study on the contribution of completeness-aware prompt tokens.
@@
w/o $\mathbf{p}^{d}$ 27.35 0.833 16.04 27.02 0.824 17.30 26.73 0.814 19.62 25.82 0.819 16.34 24.53 0.803 33.80
w/o $\mathbf{p}^{p}$ 26.92 0.829 18.46 26.43 0.819 20.23 26.06 0.810 25.13 25.17 0.812 20.08 24.39 0.798 36.54
w/o $\mathbf{p}^{s}$ 27.56 0.835 15.26 27.22 0.827 16.37 26.90 0.815 17.58 26.27 0.828 16.79 25.08 0.813 29.83
w/o Prompts 25.92 0.823 25.69 25.06 0.807 32.17 24.83 0.802 37.97 24.70 0.797 23.29 23.56 0.778 42.17
w/ Mask Codes 27.18 0.831 17.42 26.82 0.816 20.06 26.18 0.809 24.59 26.15 0.823 18.15 24.65 0.802 35.86
CoPeDiT 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
BraTS
[…]
MMT [24] 25.19 0.824 24.53
+ Prompts (ours) 25.68 (+0.49) 0.826 (+0.002) 22.07 (–2.46)
Hyper-GAE [53] 24.65 0.813 28.97
+ Prompts (ours) 25.26 (+0.61) 0.822 (+0.009) 24.26 (–4.71)
M2DN [29] 26.45 0.830 21.29
+ Prompts (ours) 27.23 (+0.78) 0.837 (+0.007) 17.06 (–4.23)
4.1 Experimental Setup
Datasets. (i) Brain MRI Datasets. We evaluate the effectiveness of CoPeDiT on two public brain MRI datasets: BraTS 2021 [3] and IXI [5]. The BraTS 2021 dataset includes 1251 subjects with multi-modal MRI scans across four modalities: T1, T1ce, T2, and FLAIR. The IXI dataset contains 577 subjects with three MRI modalities: T1, T2, and PD. (ii) Cardiac MRI Datasets. Missing slice synthesis experiments are conducted on four cardiac MRI datasets: UK Biobank (UKBB) [35], MESA [57], ACDC [4], and MSCMR [59]. The model is trained on the combined dataset including all four sources with 32,248 MRI volumes in total, while performance comparisons are conducted on the UKBB dataset. We randomly select 80% of the data for training and use the remaining 20% as the test set. Please refer to the supplementary material for more dataset details.
Implementation Details. The compression rate of CoPeVAE is set to $(8,8,8)$. For hyperparameters, the dimension of each prompt token is set to 512, $\tau$ and $\lambda$ are set to 0.2 and 1e-2, respectively. The model is trained with a global batch size of 8 for CoPeVAE-B and 64 for CoPeVAE-C, using a learning rate of 1e-4. Regarding MDiT3D, we set the number of time steps to 500 with linearly scaled noise scheduling. The model is trained for 100k steps with a global batch size of 32 and learning rate of 5e-5. All training is conducted on four NVIDIA A100 GPUs. More details are provided in the supplementary material.
4.2 Performance Comparison
Quantitative Results. We compare our method against six recent SOTA models for 3D brain and cardiac MRI synthesis, including three GAN-based approaches [43, 24, 53] and three advanced diffusion model-based methods [39, 58, 29]. We report results by reimplementing all baselines from scratch under identical experimental settings to ensure fair comparisons. Performance is evaluated using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Fréchet Inception Distance (FID) [15] that uses Inception-V3 [46] as the feature extractor. Quantitative results on the three datasets are presented in Tables 1 and  2, where the second row indicates the number of missing modalities or slices $c$, respectively. The results show that CoPeDiT consistently outperforms all baselines across all missing configurations in both synthesis tasks, showcasing superior generalizability. Notably, the performance gains are more pronounced in scenarios with a larger number of missing modalities/slices, highlighting the robustness of our completeness-aware prompts in handling complex missing cases. Besides, CoPeDiT achieves significantly lower FID, indicating that the generated MRIs are more anatomically coherent and texture-preserving, with feature distributions that are closer to real clinical data, thus enhancing perceptual realism and diagnostic plausibility.
Qualitative Results. As depicted in Fig. 4 and 5, our model generates synthetic MRIs that exhibit the highest visual similarity to the ground truth images, particularly in accurately capturing tumor regions. Our CoPeDiT excels at preserving subtle textural details and modeling complex anatomical structures within brain tissues, justifying our motivation that incorporating completeness perception leads to improved anatomical consistency and realism.
Table 6: Ablation study on the choice of prompt injection positions within the MDiT3D blocks. ‘Spatial / Planar’ indicates that prompt tokens are injected into the spatial and planar blocks for brain and cardiac tasks, respectively. ‘Both’ indicates that prompt tokens are injected into both the spatial and modal blocks for brain tasks, and into both the planar and spatial blocks for cardiac tasks. ‘Modal / Spatial’ indicates that prompt tokens are injected into the modal and spatial blocks, respectively.
BraTS UKBB
1 2 3 8 24
[…]
Spatial / Planar 27.47 0.834 15.41 27.20 0.825 16.83 27.03 0.817 17.45 26.29 0.827 16.82 25.22 0.814 29.24
Both 27.76 0.836 13.94 27.54 0.827 15.48 27.29 0.818 16.77 26.33 0.827 16.39 25.28 0.815 28.45
Modal / Spatial 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84
@@ 4.2 Performance Comparison
UNet [39] 25.61 0.825 23.56 25.20 0.816 28.37 24.78 0.805 36.70 24.46 0.800 22.34 23.01 0.778 41.59
DiT [34] 26.78 0.835 19.72 26.14 0.823 21.84 25.74 0.808 24.47 25.89 0.825 16.28 24.78 0.799 36.10
DiT-3D [30] 26.84 0.835 19.23 26.23 0.822 21.95 25.90 0.810 26.08 26.08 0.825 16.37 24.83 0.806 33.73
MDiT3D 28.26 0.842 12.67 28.13 0.831 13.25 27.91 0.822 14.89 26.42 0.831 15.53 25.39 0.817 25.84

### v2

@@ 3.2 Stage I: Completeness Perception Tokenizer
LDM Rombach et al. (2022) 23.84 $\pm$1.52 0.805 $\pm$0.019 36.47 740.26 23.12 $\pm$1.64 0.798 $\pm$0.021 45.93 897.83 22.65 $\pm$1.77 0.791 $\pm$0.025 52.50 1161.21
ControlNet Zhang et al. (2023) 23.98 $\pm$1.49 0.808 $\pm$0.018 34.09 806.82 23.34 $\pm$1.60 0.801 $\pm$0.020 41.28 986.30 22.85 $\pm$1.68 0.795 $\pm$0.022 48.07 1074.23
M2DN Meng et al. (2024) 26.45 $\pm$1.36 0.830 $\pm$0.016 21.29 376.53 25.87 $\pm$1.48 0.820 $\pm$0.017 27.36 467.66 25.21 $\pm$1.59 0.809 $\pm$0.024 32.40 553.16
DiffM4RI Ye et al. (2026) 26.07 $\pm$1.57 0.824 $\pm$0.024 25.03 449.17 25.49 $\pm$1.41 0.813 $\pm$0.025 32.59 591.35 25.08 $\pm$1.75 0.806 $\pm$0.029 35.91 695.06
CoPeDiT 28.26 $\pm$1.24 0.842 $\pm$0.019 12.67 254.71 28.13 $\pm$1.49 0.831 $\pm$0.021 13.25 287.58 27.91 $\pm$1.41 0.822 $\pm$0.023 14.89 323.19
Table 2: Quantitative results for multi-modal brain MRI synthesis on the IXI dataset. The numbers in the first row denote the number of missing modalities.
1 2
[…]
GAN-based Methods
MMGAN Sharma & Hamarneh (2020) 22.29 $\pm$1.35 0.684 $\pm$0.015 70.91 1447.83 21.13 $\pm$1.49 0.668 $\pm$0.019 93.57 1787.39
MMT Liu et al. (2023) 22.64 $\pm$1.49 0.698 $\pm$0.013 53.60 1329.25 21.82 $\pm$1.60 0.687 $\pm$0.019 72.24 1562.44
Hyper-GAE Yang et al. (2023) 22.12 $\pm$1.33 0.682 $\pm$0.017 72.62 1520.49 20.91 $\pm$1.45 0.662 $\pm$0.021 98.79 1712.90
Diffusion Model-based Methods
LDM Rombach et al. (2022) 21.36 $\pm$1.28 0.679 $\pm$0.016 86.62 1884.64 20.94 $\pm$1.53 0.654 $\pm$0.020 112.57 2314.39
ControlNet Zhang et al. (2023) 21.83 $\pm$1.51 0.681 $\pm$0.015 81.26 1971.40 21.07 $\pm$1.50 0.661 $\pm$0.019 103.77 2386.04
M2DN Meng et al. (2024) 23.47 $\pm$1.43 0.715 $\pm$0.014 42.52 845.29 22.81 $\pm$1.56 0.702 $\pm$0.018 55.64 1078.36
DiffM4RI Ye et al. (2026) 23.71 $\pm$1.26 0.720 $\pm$0.018 39.87 715.63 22.87 $\pm$1.68 0.707 $\pm$0.022 51.93 964.28
CoPeDiT 24.34 $\pm$1.21 0.732 $\pm$0.016 25.84 569.22 23.92 $\pm$1.53 0.721 $\pm$0.020 32.53 718.54
@@ 3.3 Stage II: 3D MRI Diffusion Transformer
We propose MDiT3D (see Fig. 3), a task-specific diffusion transformer extending DiT Peebles & Xie (2023) for volumetric MRI synthesis. MDiT3D is conditionally guided by the available latents $\mathbf{z}^{s}$ alongside the concatenated completeness-aware prompts $\mathbf{p}=\mathbf{p}^{d}\,\|\,\mathbf{p}^{p}\,\|\,\mathbf{p}^{s}$. During inference, these prompts are autonomously extracted by the frozen CoPeVAE to provide semantic guidance: $\mathbf{p}^{d}$ calibrates global severity (how many), $\mathbf{p}^{p}$ localizes missing regions (where), and $\mathbf{p}^{s}$ supplies fine-grained textural priors (what). To effectively adapt to 3D medical data, we introduce customized operations (using MDiT3D-B as a representative example). Specifically, we apply a 3D patchify operator to project inputs into tokens $\mathbf{h}\in\mathbb{R}^{m\times T\times d}$ Mo et al. (2023), where $m$ is the number of modalities/slices and $d$ is the embedding dimension. Unlike standard 2D implementations Dosovitskiy et al. (2020), we incorporate 3D frequency-based sine-cosine positional embeddings (3D PE) to preserve precise spatial relationships.
Alternating Blocks & Prompt Injection. We design two distinct alternating block architectures tailored to the specific characteristics of brain and cardiac MRIs. For the multi-modal brain task, we alternate between Spatial Blocks (to capture 3D spatial context) and Modal Blocks (to model inter-modal relationships). For the volumetric cardiac task, we alternate between Planar Blocks (for intra-slice features) and Spatial Blocks (for through-plane continuity). To compute attention along the appropriate axes, the latent tokens are dynamically reshaped before entering each block (e.g., aligning the modal dimension to facilitate inter-modal interaction, and then reshaped back for spatial processing). Furthermore, to maximize the efficacy of conditional guidance, prompts are injected via adaptive layer normalization (adaLN) Peebles & Xie (2023) exclusively into the blocks that explicitly model the task’s primary dependency. Specifically, prompt tokens are injected only into the Modal Blocks for brain MRI (addressing missing modalities) and the Spatial Blocks for cardiac MRI (addressing missing slices). This targeted injection avoids overwhelming the network and ensures the conditioning signals are both physically meaningful and highly informative.
Joint Reconstruction & Synthesis. During diffusion, rather than filling missing modalities or slices with zeros or learned tokens, we only add noise to the missing sections, keeping the available latents unperturbed to provide rich contextual guidance. Following Meng et al. (2024), MDiT3D is optimized via an $\mathbf{x}_{0}$-prediction loss:
@@
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{\mathbf{z}_{0},\epsilon,t}\left[\left\|\mathbf{z}_{0}-f_{\theta}(\mathbf{z}_{t},t,\mathbf{p})\right\|_{2}^{2}\right],$ (5)
@@
where $\mathbf{z}_{0}=\mathbf{z}^{s}\,\|\,\mathbf{z}^{c}$ is the clean target, $\mathbf{z}_{t}=\mathbf{z}^{s}\,\|\,\mathbf{z}_{t}^{c}$ is the partially noised input at timestep $t$, $\mathbf{p}$ represents the prompt tokens, and $f_{\theta}$ denotes MDiT3D.
@@
Table 3: Quantitative results for Cardiac MRI synthesis on the UKBB dataset. The numbers in the first row denote the length of missing slices.
@@
MMGAN Sharma & Hamarneh (2020) 25.81 $\pm$0.86 0.815 $\pm$0.012 16.68 392.48 24.35 $\pm$0.84 0.793 $\pm$0.009 27.52 604.37 23.06 $\pm$1.08 0.776 $\pm$0.018 45.88 767.92
MMT Liu et al. (2023) 26.02 $\pm$0.82 0.824 $\pm$0.009 15.90 429.72 24.73 $\pm$0.87 0.809 $\pm$0.010 23.75 570.29 24.12 $\pm$1.06 0.794 $\pm$0.017 37.39 708.68
Hyper-GAE Yang et al. (2023) 25.23 $\pm$0.89 0.810 $\pm$0.012 19.02 517.55 23.66 $\pm$0.94 0.789 $\pm$0.013 31.84 648.35 22.70 $\pm$1.14 0.771 $\pm$0.019 48.27 845.14
@@
LDM Rombach et al. (2022) 24.17 $\pm$0.96 0.795 $\pm$0.012 24.61 704.91 23.04 $\pm$1.03 0.778 $\pm$0.010 44.93 830.73 22.19 $\pm$1.21 0.761 $\pm$0.018 60.02 910.15
ControlNet Zhang et al. (2023) 24.62 $\pm$0.91 0.801 $\pm$0.012 22.47 698.63 23.30 $\pm$0.98 0.784 $\pm$0.014 37.51 821.40 22.46 $\pm$1.16 0.765 $\pm$0.020 54.93 922.06
M2DN Meng et al. (2024) 25.48 $\pm$0.79 0.814 $\pm$0.011 17.72 453.03 24.62 $\pm$0.87 0.803 $\pm$0.011 24.15 597.78 24.03 $\pm$0.99 0.780 $\pm$0.021 40.08 823.70
DiffM4RI Ye et al. (2026) 25.19 $\pm$0.89 0.808 $\pm$0.014 19.16 495.23 24.78 $\pm$0.80 0.807 $\pm$0.015 22.96 548.44 24.27 $\pm$1.06 0.791 $\pm$0.018 35.27 674.12
CoPeDiT 26.42 $\pm$0.81 0.831 $\pm$0.013 15.53 318.62 26.07 $\pm$0.74 0.826 $\pm$0.013 18.21 382.04 25.39 $\pm$0.86 0.817 $\pm$0.016 25.84 490.57
4 Experiments and Results
4.1 Experimental Setup
Datasets. (i) Brain MRI Datasets. We evaluate the effectiveness of CoPeDiT on two public brain MRI datasets: BraTS 2021 Baid et al. (2021) and IXI Brain Development Project (2025). The BraTS 2021 dataset includes 1251 subjects with multi-modal MRI scans across four modalities: T1, T1ce, T2, and FLAIR. The IXI dataset contains 577 subjects with three MRI modalities: T1, T2, and PD. (ii) Cardiac MRI Datasets. Missing slice synthesis experiments are conducted on four cardiac MRI datasets: UK Biobank (UKBB) Petersen et al. (2016), MESA Zhang et al. (2018), ACDC Bernard et al. (2018), and MSCMR Zhuang et al. (2022). The model is trained on the combined dataset including all four sources with 32,248 MRI volumes in total, while performance comparisons are conducted on the UKBB dataset. We randomly select 80% of the data for training and use the remaining 20% as the test set. Please refer to Appendix A for more details.
Implementation Details. The compression rate of CoPeVAE is set to $(8,8,8)$. For hyperparameters, the dimension of each prompt token is set to 512, $\tau$ and $\lambda$ are set to 0.2 and 1e-2, respectively. The model is trained with a global batch size of 8 for CoPeVAE-B and 64 for CoPeVAE-C, using a learning rate of 1e-4. Regarding MDiT3D, we set the number of time steps to 500 with linearly scaled noise scheduling. The model is trained for 100k steps with a global batch size of 32 for MDiT3D-B and 64 for MDiT3D-C, with learning rate of 5e-5. All training is conducted on four NVIDIA A100 GPUs. More details are provided in Appendix B.
Figure 4: Qualitative results on the BraTS dataset. The visual results on the IXI dataset are provided in Appendix C.1.
Figure 5: Qualitative results on the UKBB dataset. The top and bottom results correspond to the first and last missing slices within a given volume, respectively.
Table 4: Ablation study on the contribution of completeness-aware prompts.
BraTS UKBB
1 2 3 8 24
[…]
w/o $\mathbf{p}^{d}$ 27.35 0.833 16.04 335.08 27.02 0.824 17.30 475.32 26.73 0.814 19.62 556.79 25.82 0.819 16.34 427.06 24.53 0.803 33.80 610.46
w/o $\mathbf{p}^{p}$ 26.92 0.829 18.46 372.36 26.43 0.819 20.23 460.45 26.06 0.810 25.13 572.03 25.17 0.812 20.08 534.80 24.39 0.798 36.54 697.51
w/o $\mathbf{p}^{s}$ 27.56 0.835 15.26 341.70 27.22 0.827 16.37 395.43 26.90 0.815 17.58 468.23 26.27 0.828 16.79 388.46 25.08 0.813 29.83 578.13
w/o Prompts 25.92 0.823 25.69 418.12 25.06 0.807 32.17 556.92 24.83 0.802 37.97 748.25 24.70 0.797 23.29 721.35 23.56 0.778 42.17 830.72
w/ Mask Codes 27.18 0.831 17.42 356.09 26.82 0.816 20.06 439.13 26.18 0.809 24.59 543.82 26.15 0.823 18.15 409.47 24.65 0.802 35.86 612.73
CoPeDiT 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57
[…]
MMT Liu et al. (2023) 25.19 0.824 24.53 527.06
+ Prompts (ours) 25.68 (+0.49) 0.826 (+0.002) 22.07 (–2.46) 416.34 (–110.72)
Hyper-GAE Yang et al. (2023) 24.65 0.813 28.97 609.65
+ Prompts (ours) 25.26 (+0.61) 0.822 (+0.009) 24.26 (–4.71) 523.75 (–85.90)
M2DN Meng et al. (2024) 26.45 0.830 21.29 376.53
+ Prompts (ours) 27.23 (+0.78) 0.837 (+0.007) 17.06 (–4.23) 308.68 (–67.85)
Table 6: Ablation study of each pretext task’s contribution to CoPeVAE’s reconstruction capacity. ’Cardiac’ refers to the combined dataset.
@@ 4.1 Experimental Setup
4.2 Performance Comparison
We compare our method against seven SOTA baselines, including three GAN-based approaches Sharma & Hamarneh (2020); Liu et al. (2023); Yang et al. (2023) and four diffusion-based models Rombach et al. (2022); Zhang et al. (2023); Meng et al. (2024); Ye et al. (2026), all reimplemented under identical settings for fair comparison. Performance is evaluated using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), Fréchet Inception Distance (FID) Heusel et al. (2017), and Fréchet Video Distance (FVD) Unterthiner et al. (2018) for assessing 3D spatial consistency.
Quantitative Results. Quantitative results on the three datasets are presented in Tables 12 and 3, respectively. CoPeDiT consistently outperforms all baselines across all missing configurations in both synthesis tasks. Notably, these performance gains widen in scenarios with a higher number of missing modalities or slices, such as maintaining a high 27.91 PSNR even with three missing modalities. This highlights the robustness of our completeness-aware prompts in complex cases. Moreover, CoPeDiT achieves substantially lower FID and FVD scores (e.g., an FVD of 490.57 for 24 missing cardiac slices ). This indicates that the generated MRIs are not only anatomically coherent and texture-preserving, but exhibit superior 3D spatial consistency and structural continuity, enhancing perceptual realism and diagnostic plausibility.
Qualitative Results. As depicted in Figs. 4 and 5, our model generates synthetic MRIs that exhibit the highest visual similarity to the ground truth images, particularly in accurately capturing tumor regions. Our CoPeDiT excels at preserving subtle textural details and modeling complex anatomical structures within brain tissues, justifying our motivation that incorporating completeness perception leads to improved anatomical consistency and realism.
Table 7: Ablation study on the choice of prompt injection positions within the MDiT3D blocks. The labels denote the target blocks for brain and cardiac tasks, respectively (e.g., "Modal / Spatial" means injecting prompts exclusively into the modal blocks for brain MRIs and spatial blocks for cardiac MRIs). "Both" indicates injection into both types of blocks for the respective tasks.
@@ 4.2 Performance Comparison
Spatial / Planar 27.47 0.834 15.41 317.62 27.20 0.825 16.83 484.39 27.03 0.817 17.45 592.17 26.29 0.827 16.82 405.72 25.22 0.814 29.24 604.32
Both 27.76 0.836 13.94 294.37 27.54 0.827 15.48 367.20 27.29 0.818 16.77 434.83 26.33 0.827 16.39 414.37 25.28 0.815 28.45 587.13
Modal / Spatial 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57
@@
UNet Rombach et al. (2022) 25.61 0.825 23.56 433.55 25.20 0.816 28.37 545.89 24.78 0.805 36.70 707.83 24.46 0.800 22.34 689.53 23.01 0.778 41.59 871.64
DiT Peebles & Xie (2023) 26.78 0.835 19.72 392.09 26.14 0.823 21.84 498.13 25.74 0.808 24.47 611.26 25.89 0.825 16.28 403.83 24.78 0.799 36.10 592.72
DiT-3D Mo et al. (2023) 26.84 0.835 19.23 359.83 26.23 0.822 21.95 487.12 25.90 0.810 26.08 593.42 26.08 0.825 16.37 397.71 24.83 0.806 33.73 610.89
MDiT3D 28.26 0.842 12.67 254.71 28.13 0.831 13.25 287.58 27.91 0.822 14.89 323.19 26.42 0.831 15.53 318.62 25.39 0.817 25.84 490.57